OpenStudy (anonymous):

in a class, students study either history or geography, but not both. If the number of students studying Geography is 50% more than those studying History, what is the percentage of students studying history? A. 25% B. 33 1/3% C. 40% D. 60% E. 66 2/3%

OpenStudy (anonymous):

geography majority maybe 51% and history minority maybe 49%.

OpenStudy (anonymous):

but it sais its 50%

OpenStudy (anonymous):

@Navycoolguy yes, i agree with you

OpenStudy (anonymous):

so it cant be 51 pecent because it sais 50% more so how much is 50%

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

half the amount of students right?

OpenStudy (anonymous):

@Navycoolguy i don't think so

OpenStudy (anonymous):

hmm let me read it one more time

OpenStudy (anonymous):

50% more than history

OpenStudy (anonymous):

wait a second its saying the the amout of students studying history is more then half the amout studying geography

OpenStudy (anonymous):

@fuli yes

OpenStudy (anonymous):

@Navycoolguy 50%more , it may not all students study ONLY history and geog

OpenStudy (anonymous):

thats true

OpenStudy (anonymous):

@genius12 @Callisto

OpenStudy (callisto):

Lol I did this question a week ago... Let the no. of students in the class be C, no of students who study History be H and no of students who study Geography be G, where C = G+H Since number of students studying Geography is 50% more than those studying History, G = (1+50%) H = 1.5H Percentage of students studying history = H/C *100% = ...

OpenStudy (anonymous):

40%?

OpenStudy (anonymous):

I think the question would be easier if there were some numbers involved not just percents

OpenStudy (callisto):

You got it.

OpenStudy (anonymous):

thanks